: Devise a strategy to derive an expression of Electric Field at a point R due to an Electric Dipole without using the expression of Electric Scalar Potential. You can use either Coulumb’s or Gauss’s Law.
1
Expert's answer
2020-11-20T09:35:37-0500

As point P is very far from the dipole, given dipole can be considered as a point dipole.

From Coulumb’s Law for a point dipole electric field on axial line is

"E_a = \frac{2kp}{R^3}"

p = qd

Electric field on equatorial line is

"E_e = -\frac{kp}{R^3}"

"E_2 = \frac{2kpcosθ}{R^3}"

"E_1 = -\frac{kpsinθ}{R^3}"

The net field at P is

"E = \sqrt{E_1^2 + E_2^2} = \frac{kp}{R^3}\sqrt{sin^2θ + 4cos^2θ} = \frac{kp}{R^3}\sqrt{1 + 3cos^2θ}"